Two planes leave O'Hare Airport in Chicago a the same time. One flies directly east at 540 miles per hour. The other flies directly west at 720 miles per hour. How long will it take them to be 1890 miles apart? Use the five step method to solve.
Answer by checkley75(3666)

D=RT
1890=540x+720x
1890=1260x
x=1890/1260
x=1.5 HOURS THEY WILL BE 1890 MILES APART
PROOF:
1890=540*1.5+720*1.5
1890=810+1080
1890=1890